OpenStudy (carlyleukhardt):

The perimeter of a rectangle is 34 inches. If the length of the rectangle is 11 inches, which equation could be used to find the width, x? 11(x + 2) = 34 2(x + 11) = 34 x + 11 = 34 x + 2(11) = 34

OpenStudy (anonymous):



OpenStudy (carlyleukhardt):

Is it c?

OpenStudy (anonymous):

you got \[x+x+11+11=34\] or since you are given silly choices go with \[3(x+11)=34\]

OpenStudy (anonymous):

no it is not C the perimeter is all of the sides

OpenStudy (carlyleukhardt):

There isn't an answer choice for that?

OpenStudy (anonymous):

lol i made a typo \[2(x+11)=34\]

OpenStudy (carlyleukhardt):

Ahhh thank you <3 :D xD